OpenStudy (anonymous):

Please help c: Solve 2sqrt x+7 – 10 = –8 x = –6 x = –4 x = 32 x = 74

OpenStudy (luigi0210):

\[2\sqrt{x+7}-10=-8\]

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

Which quadratic equation has a vertex of (0, 5)? y = (x – 5)2 y = (x + 5)2 y = x2 – 5 y = x2 + 5

OpenStudy (anonymous):

@Nurali

OpenStudy (nurali):

The vertex form of a quadratic equation is: y = a(x-h)²+k where -(h) is the x-coordinate and (k) is the y-coordinate. So, if you plug this point into the formula, you get: y = a(x-0)²+5 Since the value of (a) is unknown and irrelevant to the question, it can be assumed to be 1. simplified: y = x²+5

OpenStudy (luigi0210):

first add 10\[2\sqrt{x+7}=2\] \[\sqrt{x+7}=1\] \[(x+7)=1^2\] x=-6

OpenStudy (anonymous):

Makes sense!

OpenStudy (anonymous):

Thank u

OpenStudy (anonymous):

lol "makes sense" translated = ty for doing the work and giving me the answer..now i don't have to learn a thing and i can keep smiling and asking for more answers

OpenStudy (anonymous):

:D LOL

OpenStudy (anonymous):

Im definitely smiling.

OpenStudy (anonymous):

rofl

OpenStudy (luigi0210):

Yay, we made someone smile!